OpenStudy (anonymous):

the sum of two numbers is 165 and thier diferrence is 29, fid the numbers?

OpenStudy (anonymous):

x+y = 165 x-y = 29 now solve the system of equation using substitution :)

OpenStudy (anonymous):

show me how

OpenStudy (anonymous):

for the first equation, isolate y. so y = 165 - x then substitute the first equation into the 2nd equation x-(165-x) = 29 simplify 2x-165 = 29 2x = 194 x = 97. then substitute x into one of the equation, i'll use equation 2 97-y = 29 isolate the variable -y = -68 y = 68. the numbers are 68 and 97

OpenStudy (anonymous):

x+y = 165 x-y = 29 2x=194 x=97 97+y=165 y=68

OpenStudy (anonymous):

please award good answer so people know ur question is answered.